• Monday 18/09 meeting
    • Rolandas reported on temporary fix for RB /tmp filling up. Scripts are reporting very verbose mode and are hardwired.
    • James noted scheduled start of CMS activity and need to closely monitor for gLite RB problems.

    • broadcasts - "Tape Lib intervention 27/9","Short notice on a LSF restart"

  • Tuesday meeting
    • nothing of note

    • broadcasts - "Short SFT interruption"

  • Wednesday meeting
    • Remedy 367422 - glite-wms-check-daemons. Closed on open savannah ticket 19929 for yaim config to change.

    • CCSR - intermittent (~each 2 weeks) site firewall blocks needing router reboot. Will be announced by MOD.

    • broadcasts - "CERN FTS Web Server intermittent problems"

  • Thursday meeting
    • nothing of note
    • WLCG coordination meeting
      • nothing of note

    • broadcasts - "ce102.cern.ch problems due to high load"

  • Friday meeting
    • nothing of note

-- Main.ineilson - 18 Sep 2006

Edit | Attach | Watch | Print version | History: r6 < r5 < r4 < r3 < r2 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r6 - 2006-09-22 - IanRobertNeilson
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2019 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback